# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 2.1 A brief review of MFG and fictitious play
$\quad~{}~{}\rho(x,0)=\rho_{0}(x),x\in\Omega,$
$\quad~{}~{}\nabla\rho(x,t)\cdot\mathbf{n}=0,\quad~{}~{}m(x,t)\cdot\mathbf{n}=0,(x,t)\in\partial\Omega\times(0,T).$
Here, $m=\rho v$ is the momentum and when $\rho=0$ the quotient is defined by convention
$L\left(x,\frac{m}{\rho}\right)=\begin{cases}0,&\text{if }m=0\\ +\infty,&\text{otherwise}.\end{cases}$ (9)
When $L$ is convex, $f$ and $f_{T}$ are monotone, (8) is an optimization problem with a convex objective and linear constraint. The minimizer is of the form $(\rho,-\rho D_{p}H(x,\nabla\phi))$ where $(\rho,\phi)$ solves (2). Some first-order optimization algorithms such as proximal gradient [51], primal-dual [44] and augmented Lagrangian [9, 10] have been adapted to solve (8). While introducing $m=\rho v$ gives the variational formulation many nice properties, it makes the Lipschitz constant of the objective function extremely large when $\rho$ is close to 0 and causes algorithms such as proximal gradient [51] to be unstable.
2.2 Convergence and acceleration of fictitious play
Given the fictitious play algorithm (Algorithm 1), several questions arise: Does it converge? Under what assumptions does convergence occur? And if it does, what is the rate of convergence? For potential MFGs, the objective function serves as a natural criterion for studying convergence. Based on this, [19] demonstrates the convergence of fictitious play for both first-order and second-order potential games, and [41] offers a sublinear convergence of classical fictitious play through the lens of the generalized conditional gradient. [41] also shows that a linear convergence rate can be achieved when the weight $\delta_{k}$ is chosen wisely.
In this section, we aim to demonstrate the linear convergence of fictitious play for non-potential MFGs. To this end, we introduce the important concept of the best response in Section 2.2.1. With this, we reformulate the fictitious play algorithm and define the gain of the best response as a measure of the distance to the Nash Equilibrium. In Section 2.2.2, we analyze the stability of the best-response mapping, a crucial step in understanding the fictitious play. Leveraging these stability properties, we show that the gain of the best response functions is a Lyapunov function and prove the convergence of fictitious play for non-potential MFGs; details are provided in Section 2.2.3. Section 2.2.4 presents the backtracking line search for optimizing the weight $\delta_{k}$ to accelerate the algorithm. Finally, Section 2.2.5 remarks on the convergence for potential MFGs, relating the gain of the best response to the dual gap in the generalized conditional gradient (Frank-Wolfe) algorithm.
2.2.1 The best response and the equilibrium formulation
We denote the constraint set as
$C:=\{({\rho},{m}):\partial_{t}{\rho}(x,t)-\nu\Delta{\rho}(x,t)+\nabla\cdot{m}(x,t)=0,~{}{\rho}(x,0)=\rho_{0}(x),\rho_{t}\in\mathcal{P}(\Omega)\cap L^{2}(\Omega)\}.$ (10)
$C$ is convex because the constraint is linear. We define the cost of $(\tilde{\rho},\tilde{m})$ given $\rho$ as
$J(\tilde{\rho},\tilde{m};\rho):=r(\tilde{\rho},\tilde{m})+\langle f(\cdot,\rho),\tilde{\rho}\rangle_{Q_{T}}+\langle f_{T}(\cdot,\rho_{T}),\tilde{\rho}_{T}\rangle_{\Omega},$ (11)
where
$r(\tilde{\rho},\tilde{m}):=\int_{0}^{T}\int_{\Omega}\tilde{\rho}(x,t)L\left(x,\frac{\tilde{m}(x,t)}{\tilde{\rho}(x,t)}\right)\mathrm{d}x\mathrm{d}t,$
and
$\langle f(\cdot,\rho),\tilde{\rho}\rangle_{Q_{T}}=\int_{0}^{T}\int_{\Omega}f(x,\rho_{t})\mathrm{d}\tilde{\rho}_{t}(x)\mathrm{d}t,\quad\langle f_{T}(\cdot,\rho_{T}),\tilde{\rho}_{T}\rangle_{\Omega}=\int_{\Omega}f_{T}(x,\rho_{T})\mathrm{d}\tilde{\rho}_{T}(x).$
Here the dynamic cost $r(\tilde{\rho},\tilde{m})$ is convex in $(\tilde{\rho},\tilde{m})$, the interaction cost $\langle f(\cdot,\rho),\tilde{\rho}\rangle_{Q_{T}}$ and the terminal cost $\langle f_{T}(\cdot,\rho_{T}),\tilde{\rho}_{T}\rangle_{\Omega}$ are linear in $\tilde{\rho}$. The following lemma gives basic properties of the optimal control problem
$\inf_{\tilde{\rho},\tilde{m}\in C}J(\tilde{\rho},\tilde{m};\rho).$
Lemma 2.4.
For any $\rho\in\mathcal{P}(\Omega)$, the following statements are true.
  1. 1.
    If $f_{T}(\cdot,\rho_{T})$ is Lipschitz continuous, then there exists a unique $(\hat{\rho},\hat{\phi})$ solving the decoupled system
    $-\partial_{t}\hat{\phi}(x,t)-\nu\Delta\hat{\phi}(x,t)+H(x,\nabla\hat{\phi}(x,t))=f(x,\rho(\cdot,t)),$ $(x,t)\in Q_{T},$ (12)
    $\hat{\phi}(x,T)=f_{T}(x,\rho(T)),\quad\hat{\rho}(x,0)=\rho_{0}(x),$ $x\in\Omega.$
    If in addition, $\nu>0$, then $\hat{\rho}(x,t)>0$ for any $x\in\Omega,t\in(0,T]$.
  2. 2.
    There exists $(\hat{\rho},\hat{m})$ such that $(\hat{\rho},\hat{m})\in\operatorname*{Argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)$. If $(\hat{\rho},\hat{m})\in\operatorname*{Argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)$ and $\nu>0$, then $\hat{m}=-\hat{\rho}D_{p}H(x,\nabla\hat{\phi})$ where $(\hat{\rho},\hat{\phi})$ solves (12) and $(\hat{\rho},\hat{m})$ is the unique minimizer.
@@ 2.2.1 The best response and the equilibrium formulation
The first statement is from the classical results of the HJB equation and FP equation. The second statement can be proved by deriving the Euler-Lagrangian equation of $\min_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)$.
$-\partial_{t}\hat{\phi}(x,t)-\nu\Delta\hat{\phi}(x,t)+\left(\frac{\hat{m}}{\hat{\rho}}D_{v}L\left(x,\frac{\hat{m}}{\hat{\rho}}\right)-L\left(x,\frac{\hat{m}}{\hat{\rho}}\right)\right)=f(x,\rho(\cdot,t)),$ $(x,t)\in Q_{T},$ (13)
$\partial_{t}\hat{\rho}(x,t)-\nu\Delta\hat{\rho}(x,t)+\nabla\cdot\hat{m}(x,t)=0,$ $(x,t)\in Q_{T},$
$D_{v}L\left(x,\frac{\hat{m}}{\hat{\rho}}\right)+\nabla\hat{\phi}=0,$ $(x,t)\in Q_{T},$
$\hat{\phi}(x,T)=f_{T}(x,\rho(T)),\quad\hat{\rho}(x,0)=\rho_{0}(x),$ $x\in\Omega.$
Since $L(x,v)$ is convex in $v$ and $H(x,p)$ is the convex conjugate of $L$ with respect to $v$, the third equation of (13) gives $\hat{m}=-\hat{\rho}D_{p}H(x,\nabla\hat{\phi})$ and $\frac{\hat{m}}{\hat{\rho}}D_{v}L\left(x,\frac{\hat{m}}{\hat{\rho}}\right)-L\left(x,\frac{\hat{m}}{\hat{\rho}}\right)=H(x,\nabla\hat{\phi}).$
@@
For any $\rho\in\mathcal{P}(\Omega)$, $(\hat{\rho},\hat{m})$ is called a best response of $\rho$ if it solves the minimization problem
$(\hat{\rho},\hat{m})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho).$ (14)
And we define the gain of the best response as the gap between $J$ evaluated at $\rho$ and its best response, i.e.
$g(\rho,m):=J(\rho,m;\rho)-\min_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho).$ (15)
@@
Our definition of “best response” is related to but different from the “best reply strategy” in game theory. The relation between MFG and best reply strategy is studied in [22, 21]. Additionally, our concept of “the gain of the best response” aligns with the notion of “exploitability” in related literature. For instance, [30] introduces this concept for first-order monotone MFGs and uses it to prove convergence. Similarly, [25] employs exploitability to demonstrate first-order convergence of fictitious play in continuous-time discrete MFGs.
While we write $g$ as a functional of both $\rho$ and $m$, $m$ is an intermediate step to obtain $\rho$ and does not play a key role in later analysis. A mean-field Nash Equilibrium is the fixed point of the best response mapping, or equivalently the root of the gain of the best response, as detailed in the following Lemma.
@@
If $(\rho^{*},m^{*})\in C$ and $\rho^{*}(x,t)>0$ for any $(x,t)\in Q_{T}$, then the following statements are equivalent,
@@
    $(\rho^{*},m^{*})=(\rho^{*},-\rho^{*}D_{p}H(x,\nabla\phi^{*}))$ where $(\rho^{*},\phi^{*})$ solves (2)
@@
$1\Leftrightarrow 2$ is from Lemma 2.4. $2\Leftrightarrow 3$ is by definition. ∎
In the rest of this paper, we refer (2) the PDE formulation, (8) the variational formulation and
$(\rho,m)=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)\qquad\text{ or }\qquad g(\rho,m)=0$
the equilibrium formulation of a MFG.
The best response is also the key concept to analyze the convergence of fictitious play for non-potential MFGs. Lemma 2.4 shows that under mild conditions, the fictitious play (Algorithm 1) is equivalent to iteratively solving for the best response and computing the weighted average of the replies. To be precise, it is equivalent to the following scheme,
and the $\hat{\phi}^{(k)}$ in Algorithm 1 is the Lagrangian multiplier of (16).
2.2.2 The stability of the best response
This section studies the stability of the best response mapping, which is the key to fictitious play. Throughout, we assume there exists a Nash Equilibrium $(\rho^{*},m^{*},\phi^{*})$, i.e. $(\rho^{*},\phi^{*})$ solves (2), $(\rho^{*},m^{*})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho^{*})$ and $m^{*}=-\rho^{*}D_{p}H(\nabla\phi^{*})$. When there is no ambiguity, we omit the input $x$ of $L,H,f$ and $f_{T}$. The following assumptions and definitions appear frequently in our analysis.
Assumption 1.
There exists $L_{f}>0$ such that $f$ is $L_{f}$-Lipschitz continuous in $\rho$ for any $x\in\Omega$.
Assumption 2.
@@ 2.2.2 The stability of the best response
Definition 2.8 (Directional strong convexity and Lipschitz smooth).
For given $(\rho^{i},m^{i})\in C,i=0,1$, denote $({\rho}^{\epsilon},{m}^{\epsilon}):=(1-\epsilon)({\rho}^{0},{m}^{0})+\epsilon({\rho}^{1},{m}^{1})$, and
$y(\epsilon):=r\left(\rho^{\epsilon},m^{\epsilon}\right),\epsilon\in[0,1].$ (18)
We define that $r$ is directional $\lambda_{r}$-strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$ if $\lambda_{r}>0$ and $y$ is $(\lambda_{r}\|\rho^{0}-\rho^{1}\|_{Q_{T}}^{2})$-strongly convex in $[0,1]$. We define that $r$ is directional $L_{r}$-Lipschitz smooth between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$ if $L_{r}>0$ and $y$ is $(L_{r}\|\rho^{0}-\rho^{1}\|_{Q_{T}}^{2})$-Lipschitz smooth in $[0,1]$.
@@
$f:\Omega\times\mathcal{P}(\Omega)$ is $\lambda_{f}$-strongly monotone if $\lambda_{f}>0$ and for any $\rho,\rho^{\prime}\in\mathcal{P}(\Omega)\cap L^{2}(\Omega)$,
$\int_{\Omega}(f(x,\rho)-f(x,\rho^{\prime}))\mathrm{d}(\rho-\rho^{\prime})(x)\geq\lambda_{f}\int_{\Omega}|\mathrm{d}(\rho-\rho^{\prime})(x)|^{2}.$

### v2

@@ 3.1 The best response and the equilibrium formulation
[…]
@@
Lemma 2.4 gives the existence of $(\hat{\rho},\hat{\phi})$. We prove that $(\hat{\rho},\hat{\phi})$ is the unique minimizer by verification. Without loss of generality, consider $(\tilde{\rho},\tilde{m})\in C$ such that $\tilde{m}=\tilde{\rho}\tilde{v}$ for some $\tilde{v}$ and $(\tilde{\rho},\tilde{m})\neq(\hat{\rho},\hat{m})$. Denote $\hat{v}:=-D_{p}H(\cdot,\nabla\hat{\phi})$ and $\hat{p}:=\nabla\hat{\phi}$. Then by definition,
$L(x,\hat{v})=-\langle\hat{p},\hat{v}\rangle-H(x,\hat{p}),$ (18)
and by strictly convexity of $H$ in the Assumption 2.3, for any $v\neq\hat{v}$,
$L(x,v)>-\langle\hat{p},v\rangle-H(x,\hat{p}).$ (19)
Therefore, we have
$R(\tilde{\rho},\tilde{m})-R(\hat{\rho},\hat{m})=$ $\int_{0}^{T}\int_{\mathbb{T}^{d}}\tilde{\rho}L(x,\tilde{v})-\hat{\rho}L(x,\hat{v})\mathrm{d}x\mathrm{d}t$ (20)
$>$ $-\left\langle\nabla\hat{\phi},\tilde{m}-\hat{m}\right\rangle_{Q_{T}}-\left\langle H(\nabla\hat{\phi}),\tilde{\rho}-\hat{\rho}\right\rangle_{Q_{T}}$
and as a result,
$J(\tilde{\rho},\tilde{m};\rho)-J(\hat{\rho},\hat{m};\rho)=$ $R(\tilde{\rho},\tilde{m})-R(\hat{\rho},\hat{m})+\langle f(\rho),\tilde{\rho}-\hat{\rho}\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\mathbb{T}^{d}}$ (21)
$>$ $\left\langle-H(\nabla\hat{\phi})+f(\rho),\tilde{\rho}-\hat{\rho}\right\rangle_{Q_{T}}-\left\langle\nabla\hat{\phi},\tilde{m}-\hat{m}\right\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\mathbb{T}^{d}}$
$=$ $\left\langle-\partial_{t}\hat{\phi}-\nu\Delta\hat{\phi},\tilde{\rho}-\hat{\rho}\right\rangle_{Q_{T}}+\left\langle\hat{\phi},\nabla\cdot(\tilde{m}-\hat{m})\right\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\mathbb{T}^{d}}$
$=$ $\left\langle\hat{\phi},(\partial_{t}-\nu\Delta)(\tilde{\rho}-\hat{\rho})+\nabla\cdot(\tilde{m}-\hat{m})\right\rangle_{Q_{T}}=0,$
where the last two equalities are due to (17). Therefore, we prove that $(\hat{\rho},\hat{m})$ is the unique minimizer.
@@
For any $(\rho,m)\in C$, we define $(\hat{\rho},\hat{m})$ as the best response of $\rho$ if it solves the minimization problem
$(\hat{\rho},\hat{m})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho),$ (22)
and define the gain of the best response as the gap between $J$ evaluated at $(\rho,m)$ and its best response:
$g(\rho,m):=J(\rho,m;\rho)-\min_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)=J(\rho,m;\rho)-J(\hat{\rho},\hat{m};\rho).$ (23)
@@
Our definition of “best response” is related to but different from the “best reply strategy” in game theory. The relation between MFG and best reply strategy is studied in degond2014large ; degond2017mfgbrs . Additionally, our concept of “the gain of the best response” aligns with the notion of “exploitability” in related literature. For instance, hadikhanloo2017anonymous introduces this concept for first-order monotone MFGs and uses it to prove convergence. Similarly, geist2022curlmfg employs exploitability to demonstrate first-order convergence of fictitious play in continuous-time discrete MFGs.
The best response is the central concept to analyze the convergence of fictitious play for non-potential MFGs. Lemma 3.1 shows that the fictitious play (Algorithm 1) is equivalent to iteratively solving for the best response and computing the weighted average of the replies. To be precise, it is equivalent to the following scheme,
and the $\hat{\phi}^{(k)}$ in Algorithm 1 is the Lagrangian multiplier of (24). In addition, a mean-field Nash Equilibrium is the fixed point of the best response mapping, or equivalently, the root of the gain of the best response, as detailed in the following Lemma.
@@
If $(\rho^{*},m^{*})\in C$ and $\rho^{*}>0$, then the following statements are equivalent,
@@
    $(\rho^{*},m^{*})=(\rho^{*},-\rho^{*}D_{p}H(\cdot,\nabla\phi^{*}))$ where $(\rho^{*},\phi^{*})$ solves (2)
@@
$1\Leftrightarrow 2$ is from Lemma 3.1, and $2\Leftrightarrow 3$ is by definition. ∎
In the rest of this paper, we refer to (2) the PDE formulation,
$\inf_{(\rho,m)\in C}J(\rho,m)$ (26)
the variational formulation and
$(\rho,m)=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho)\qquad\text{ or }\qquad g(\rho,m)=0$ (27)
the equilibrium formulation of an MFG.
3.2 Directional strong convexity and other assumptions
The following assumptions and definitions appear frequently in our analysis.
Assumption 3.5.
$f$ satisfies one of the following assumptions
  1. 1.
    $f$ is nonlocal and $\|f(\cdot,\rho)-f(\cdot,\rho^{\prime})\|_{\mathbb{T}^{d}}\leq L_{f}\|\rho-\rho^{\prime}\|_{\mathbb{T}^{d}}$ for any $\rho,\rho^{\prime}\in\mathcal{P}(\mathbb{T}^{d})$.
  2. 2.
    $f$ is local and $|f(x,\rho)-f(x,\rho^{\prime})|\leq L_{f}|\rho-\rho^{\prime}|$ for any $x\in\mathbb{T}^{d}$ and $\rho,\rho^{\prime}\in[0,+\infty)$.
Assumption 3.6.
@@ 3.2 Directional strong convexity and other assumptions
Remark 3.7 (Independent terminal cost).
The assumption that $f_{T}(x,\rho_{T})$ being independent of $\rho_{T}$ is quite common in MFG literature, see cacase2021policyitercvg ; camilli2022policyitercvgrate ; Achdou2013fdcvg ; achdou2012multigrid ; Lavigne2023gcg-ficplay ; inoue2023ficplaylsmfg for example. When this assumption holds, the terminal condition of the HJB equation is independent of $\rho_{T}$. But as long as $f(x,\rho)$ depends on $\rho$, the MFG system (2) is still coupled and the problem is not oversimplified.
@@
$f:\mathbb{T}^{d}\times\mathcal{P}(\mathbb{T}^{d})$ is $\lambda_{f}$-strongly monotone if $\lambda_{f}>0$ and for any $\rho,\rho^{\prime}\in\mathcal{P}(\mathbb{T}^{d})\cap L^{2}(\mathbb{T}^{d})$,
$\int_{\mathbb{T}^{d}}(f(x,\rho)-f(x,\rho^{\prime}))(\rho-\rho^{\prime})(x)\mathrm{d}x\geq\lambda_{f}\int_{\mathbb{T}^{d}}|\rho(x)-\rho^{\prime}(x)|^{2}\mathrm{d}x.$
Remark 3.9 (Strong monotonicity of $f$).
$f$ being strongly monotone is a common assumption in game theory and is related to the price of anarchy Gao2024finitemfg .
The next definition is the key to convergence rate estimation.
Definition 3.10 (Directional strong convexity).
For given $(\rho^{i},m^{i})\in C,i=0,1$, denote $({\rho}^{\theta},{m}^{\theta}):=(1-\theta)({\rho}^{0},{m}^{0})+\theta({\rho}^{1},{m}^{1})$, and
$y(\theta):=R\left(\rho^{\theta},m^{\theta}\right),\theta\in[0,1].$ (28)
We define that $R$ is directional $\lambda_{R}$-strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$ if $\lambda_{R}>0$ and $y$ is $(\lambda_{R}\|\rho^{0}-\rho^{1}\|_{Q_{T}}^{2})$-strongly convex in $[0,1]$.
By the strong convexity of $y$, if $\rho^{0}>0,\rho^{1}>0$ and $R$ is directional $\lambda_{R}$-strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$, then
$y^{\prime}(0)+\frac{\lambda_{R}}{2}\|\rho^{0}-\rho^{1}\|_{Q_{T}}^{2}\leq R(\rho^{1},m^{1})-R(\rho^{0},m^{0}),$
where
$y^{\prime}(\theta)=\left\langle L\left(x,\frac{m^{\theta}}{\rho^{\theta}}\right)-\frac{m^{\theta}}{\rho^{\theta}}D_{v}L\left(x,\frac{m^{\theta}}{\rho^{\theta}}\right),\rho^{1}-\rho^{0}\right\rangle_{Q_{T}}+\left\langle D_{v}L\left(x,\frac{m^{\theta}}{\rho^{\theta}}\right),m^{1}-m^{0}\right\rangle_{Q_{T}}.$
We remark that the standard definition of strong convexity requires the existence of a uniform $\lambda_{R}$ such that the function is $\lambda_{R}$-convex along any $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$. However, the dynamic cost $R$ does not satisfy this strict definition even when $L$ is strongly convex. In contrast, our proof only relies on the local existence of $\lambda_{R}$, which holds for any distinct pair $(\rho^{0},m^{0}),(\rho^{1},m^{1})\in C$ under mild conditions. To illustrate this, assume that $L$ is $C^{2}$ and uniformly $\lambda_{L}$-convex in $v$ ($\lambda_{L}>0$). For any $(\rho^{0},m^{0}),(\rho^{1},m^{1})\in C$, and $\rho^{0}>0,\rho^{1}>0$, denote
$D_{vv}^{\theta}(x,t):=D_{vv}L\left(x,\frac{m^{\theta}(x,t)}{\rho^{\theta}(x,t)}\right).$
Then
$y^{\prime\prime}(\theta)=$ $\int_{Q_{T}}\frac{1}{(\rho^{\theta})^{3}}[(\rho^{1}-\rho^{0})m^{\theta}-\rho^{\theta}(m^{1}-m^{0})]^{\top}D_{vv}^{\theta}[(\rho^{1}-\rho^{0})m^{\theta}-\rho^{\theta}(m^{1}-m^{0})]$ (29)
$\geq$ $\lambda_{L}\int_{Q_{T}}\frac{(\rho^{0}\rho^{1})^{2}}{(\rho^{\theta})^{3}}\left|\frac{m^{0}}{\rho^{0}}-\frac{m^{1}}{\rho^{1}}\right|^{2}\geq\lambda_{L}\int_{Q_{T}}\min\left\{\frac{(\rho^{0})^{2}}{\rho^{1}},\frac{(\rho^{1})^{2}}{\rho^{0}}\right\}\left|\frac{m^{0}}{\rho^{0}}-\frac{m^{1}}{\rho^{1}}\right|^{2}.$
Since $\rho^{0},\rho^{1}>0$, the lower bound is zero only if
[…]
Notice that $\rho^{\theta}$ initialize from $\rho_{0}$ and is driven by $\frac{m^{\theta}}{\rho^{\theta}}$, (30) implies $(\rho^{0},m^{0})=(\rho^{1},m^{1})$ a.e. Therefore, given the assumption that $\rho^{0}>0$ and $\rho^{1}>0$, we may assume, without loss of generality, that
$\int_{Q_{T}}\min\left\{\frac{(\rho^{0})^{2}}{\rho^{1}},\frac{(\rho^{1})^{2}}{\rho^{0}}\right\}\left|\frac{m^{0}}{\rho^{0}}-\frac{m^{1}}{\rho^{1}}\right|^{2}=\underline{c}>0.$
As a result, we obtain $y^{\prime\prime}(\theta)\geq\lambda_{L}\underline{c}>0$ for all $\theta\in[0,1]$, implying that $R$ is directionally strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$. This also highlights that the value of $\lambda_{R}$ may become small, posing numerical challenges when $\min\left\{\frac{(\rho^{0})^{2}}{\rho^{1}},\frac{(\rho^{1})^{2}}{\rho^{0}}\right\}$ is uniformly small.
On the other hand, if $f$ is $2\lambda_{f}$-strongly monotone, we can modify the definition of $R$ accordingly. Let $\lambda_{R}=\lambda_{f}$ and define the modified cost function as
$\tilde{R}(\tilde{\rho},\tilde{m}):=R(\tilde{\rho},\tilde{m})+\frac{\lambda_{R}}{2}\|\tilde{\rho}\|_{Q_{T}}^{2},\quad\tilde{f}(x,\rho):=f(x,\rho)-\lambda_{R}\rho(x),\quad\tilde{J}(\tilde{\rho},\tilde{m};\rho):=\tilde{R}(\tilde{\rho},\tilde{m})+\langle\tilde{f}(\rho),\tilde{\rho}\rangle_{Q_{T}}+\langle f_{T},\tilde{\rho}_{T}\rangle_{\mathbb{T}^{d}}.$
With this formulation, the modified cost $\tilde{R}$ is $\lambda_{R}$-convex between any $(\rho^{0},m^{0}),(\rho^{1},m^{1})\in C$, and $\tilde{f}$ is $\lambda_{f}$-monotone. The corresponding best response is given by
which leads to the modified update system:
$\hat{\phi}(x,T)=f_{T}(x,\rho(T)),\quad\hat{\rho}(x,0)=\rho_{0}(x),$ $x\in\mathbb{T}^{d}.$
When $\rho$ is close to the Nash equilibrium, the difference $\hat{\rho}-\rho$ becomes small. In this regime, the fictitious play update (3)–(4) approximately solves the modified best response problem (31), where directional strong convexity of $R$ is guaranteed. Notably, the value of $\lambda_{R}$ is constant and independent of $\rho$ and $m$.
3.3 The stability of the best response
This section examines the stability of the best response mapping. Throughout, we assume there exists a Nash Equilibrium $(\rho^{*},m^{*},\phi^{*})$, such that $(\rho^{*},m^{*})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho^{*})$ and $m^{*}=-\rho^{*}D_{p}H(\nabla\phi^{*})$.